OpenStudy (anonymous):

If a ball is thrown upwards at 20 ft/sec, the height h after t seconds is given by h = -16t2+ 20t. How many seconds does it take for the ball to reach its maximum height, and what is this height?

OpenStudy (anonymous):

A. time = 5/16 seconds height = 5 and 1/8 feet B. time = - 5/8 seconds height = 6 and 1/4 feet C. time = 5/8 seconds height = 6 and 1/4 feet D. time = - 5/16 seconds height = 5 and 1/8 feet